1. A drawer contains 10 pairs of loose, unpaired socks. Three pairs are black, 4 are gray, and 3
are brown. If you remove 2 of the socks, what is the probability that they will match?

Answer by sudhanshu_kmr(1152)


no. of black socks = 6,
gray socks = 8
brown socks = 6
For matching socks both must be either black, gray or brown.
Probability of both are black = 6C2 / 20C2
Probability of both are gray = 8C2 / 20C2
Probability of both are brown = 6C2 / 20C2
these are mutually exclusive events..
probability of socks are matching = [6C2 / 20C2] + [8C2 / 20C2] + [6C2 / 20C2]
= 15/190 + 28/190 + 15/190
= 58/190
= 29/95